OpenStudy (anonymous):

(x^2/3)*((6-x)1/3) I'm suppose to find the derivative, but I'm getting a different answer than the book HELP

OpenStudy (imstuck):

Is it this?\[f(x)=\frac{ x ^{2} }{ 3 }*(\frac{ 6-x }{ 3 })\]

OpenStudy (imstuck):

It's kind of confusing the way you typed it up there.

OpenStudy (anonymous):

no x to the 2/3 power times 6-x to the 1/3 power

OpenStudy (imstuck):

Got it! Let me work on it for a tiny bit...

OpenStudy (anonymous):

thanks

OpenStudy (imstuck):

Might sound crazy but I did some simplifying and get the function down to\[f(x)=(6x ^{2}-x ^{3})^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

let me try real quick

OpenStudy (imstuck):

I get this crazy thing...but I gave it a fightin' try...

OpenStudy (imstuck):

\[f'(x)=\frac{ 4-x }{ x(6-x)^{\frac{ 2 }{ 3 }} }\]

OpenStudy (anonymous):

yep thats it

OpenStudy (imstuck):

Sorry in advance for being so so so....right?

OpenStudy (anonymous):

can i walk you through what i did and you can tell me what i did wrong

OpenStudy (imstuck):

Are you kidding me?!

OpenStudy (imstuck):

As crazy as that looks and it's right?

OpenStudy (anonymous):

ya it sure is lol

OpenStudy (imstuck):

Lol is right!

OpenStudy (imstuck):

Ok then...I guess just simplify it first and then it's so much easier...even I could get it!

OpenStudy (anonymous):

first i derived the equation without combining. so it was.....

OpenStudy (imstuck):

it was a mess is what it was! I tried it with a multiple chain rule and it was a disaster!

OpenStudy (anonymous):

\[x^{2/3}*(6-x)^{1/3}p r i m e +(6-x)^{1/3}*x ^{2/3}p r i m e\]

OpenStudy (anonymous):

so then

OpenStudy (imstuck):

The trick here was to convert the \[(x)^{\frac{ 2 }{ 3 }}\]into\[(x ^{2})^{\frac{ 1 }{ 3 }}\]that way they both had the same index and could be multiplied out to\[(6x ^{2}-x ^{3})^{\frac{ 1 }{ 3 }}\]See? So much easier!

OpenStudy (imstuck):

Yeah, right...easier my butt...

OpenStudy (anonymous):

cool thank you!

OpenStudy (imstuck):

Wow! You are so welcome! Getting that correct just made my day! TY for the medal, also!

OpenStudy (imstuck):

is that a joint or a whistle in your mouth in your profile pic? ; )

OpenStudy (anonymous):

black and mild might as well be a joint or a blunt though